Light of wavelength 450 nm is incident normally on a diffraction grating with 300 lines per millimeter. How many orders of diffraction maxima can be obtained? =)

A. 7
B. 8
C.14
D.15

Thanks!
1
Expert's answer
2011-06-14T10:11:07-0400
D=300 lines per millimeter -> d=1/D=1/300=3.3*10^(-6) millimeter.
dsinθ=mλ ->
m=d sin θ/λ=|sin θ=1|=(3.3*10-6)/(0.45*10-6)=7,33 ≈ 7 maxima.